# Bekenstein bound on black hole entropy in non-Gaussian statistics

### v1

@@ I Introduction
Black holes provide a framework to pursue an essential link between gravitational physics and thermodynamics Bardeen:1973gs . Moreover, further studies by Bekenstein and Hawking confirmed that quantum physics plays a noteworthy role in analyzing the thermodynamic quantities of black holes Bekenstein:1973ur ; Hawking:1975vcx . In 1973, Bekenstein claimed that the black hole entropy, as a thermodynamic quantity, directly connects to its horizon area as a geometrical object Bekenstein:1972tm ; Bekenstein:1973ur . Later, Hawking approved Bekenstein’s conjecture by introducing the Bekenstein–Hawking (BH) formula $S=A/4$. Also, Hawking discovered black holes emit thermal radiation in the presence of quantum effects Hawking:1974rv ; Hawking:1975vcx . In principle, the mentioned quantum effects allow us to present an interpretation of the laws of black hole mechanics as physically corresponding to the ordinary laws of thermodynamics Wald:1999vt ; Page:2004xp ; Carlip:2014pma ; Almheiri:2020cfm . This point can be traced in the context of universal bounds on the entropy of physical systems down to the quantum level. The second law of thermodynamics tells us that the entropy of a closed physical system will grow until it reaches its maximum value. Hence, various entropy bounds have been suggested to limit the process of entropy growth in the context of a generalized second law of thermodynamics. The first attempt was made by Bekenstein using heuristic arguments of black holes in order to find an upper bound for the entropy of macroscopic bodies in terms of the total energy and radius of the smallest sphere that encloses the object Bekenstein:1980jp . Although the Bekenstein bound conjecture was obtained by studying the physics of black holes, it does not contain Newton’s constant $G$. Therefore, the Bekenstein bound is not restricted only to gravitational scenarios. Despite the quantum proofs of Bekenstein’s discovery Bekenstein:1981zz ; Bekenstein:1982zv ; Bekenstein:2000ai ; Schiffer:1989et ; Bekenstein:1990du ; Casini:2008cr , several counterexamples point out its possible shortcomings Unruh:1982ic ; Unwin:1982pg ; Page:1982fj ; Bekenstein:1983iq ; Unruh:1983ir ; Pelath:1999xt . Further studies by ’t Hooft and Susskind showed that the Bekenstein bound can not be applied to thermodynamic systems involving strong gravity tHooft:1993dmi ; Susskind:1994vu . They claimed that the entropy of our gravitational system should be bounded by area in the context of the holographic principle. Subsequently, Bousso argued that such a spacelike entropy bound is violated more broadly in many dynamical and cosmological situations. He suggested the covariant entropy bound or Bousso entropy bound by introducing a "lightsheet", defined as a region traced out by non-expanding light-rays emitted orthogonally from an arbitrary surface Bousso:1999xy . A derivation of the Bousso bound was given by Flanagan, Marolf, and Wald by considering a condition on a null hypersurface Flanagan:1999jp .
@@
In addition to the BH statistics, we might work with non-Gaussian statistics by taking into account some corrections to the BH entropy coming from non-extensivity of the horizon degrees of freedom (DOF) or quantum gravitational effects on the black hole surface area. These statistics provide a natural framework to study semiclassical effects in the context of the generalized uncertainty principle (GUP). For instance, Tsallis proposed a class of non-Gaussian entropies by considering some non-extensive effects to the Boltzmann-Gibbs statistics Tsallis:1987eu , which is parameterized by index $q$. The non-extensivity parameter $q$ refers to super-extensivity and sub-extensivity when $q<1$ and $q>1$, respectively Tsallis:2002tp . In fact, rare events occur for $10\leq q<$ 1 and frequent events occur for $q>1$ Tsallis:1998as ; NIVEN2004444 , pointing to the stretching or compressing of the entropy curve to lower or higher maximum entropy positions. The consequences of these non-extensivity effects on the thermodynamic parameters have been discussed in Abe:2001ss . In the context of the "gravity-thermodynamics" conjecture, a broad range of gravitational and cosmological scenarios have been studied in the presence of non-extensivity. Such as dark energy (DE) Barboza:2014yfe ; Nunes:2015xsa ; Zadeh:2018poj ; Ghaffari:2018wks ; Tavayef:2018xwx ; SayahianJahromi:2018irq ; Saridakis:2018unr ; Sheykhi:2018dpn ; Lymperis:2018iuz ; Sheykhi:2019bsh ; Huang:2019hex ; Aditya:2019bbk ; DAgostino:2019wko ; Mamon:2020wnh ; Mohammadi:2021wde ; Dheepika:2021fqv ; Nojiri:2022dkr , cosmic inflation Keskin:2023ngx ; Odintsov:2023vpj ; Teimoori:2023hpv , black holes deOliveira:2005jy ; Tsallis:2012js ; Komatsu:2013qia ; Mejrhit:2020dpo ; Abreu:2020wbz , gravitational waves Jizba:2024klq , the cosmic microwave background (CMB) Bernui:2005hq ; Bernui:2007wj and modified Newtonian dynamics (MOND) Abreu:2014dna ; Abreu:2018pua . A relativistic generalization of the Boltzmann-Gibbs statistics was introduced by Kaniadakis Kana , which naturally emerges from the special relativity perspective Kaniadakis:2002zz . In $\kappa$-statistics, we deal with a non-exponential distribution function parameterized by the index $-1<\kappa<1$. See its applications in holographic dark energy (HDE) Drepanou:2021jiv ; Hernandez-Almada:2021aiw ; P:2022amn , cosmic inflation Nojiri:2022dkr ; Odintsov:2023vpj ; Lambiase:2023ryq , black holes Abreu:2021avp ; Abreu:2021kwu ; Cimidiker:2023kle and other cosmological situations Luciano:2022eio . Another class of non-Gaussian statistics deals with logarithmic and power-law correction terms due to quantum gravitational effects in the physics of black holes. Considering self-gravitation and backreaction effects to the metric of the black hole navigates us to the BH logarithmic-corrected statistics described by two dimensionless and model-dependent constants $\alpha$ and $\beta$. For gravitational and cosmological applications of the logarithmic-corrected version of the BH entropy, see Refs.Jamil:2010xq ; Cai:2010zw ; Abreu:2020wbz ; Abreu:2020dyu . Recently, Barrow Barrow:2020tzx has proposed a new type of statistics by considering a fractal structure, coming from quantum gravitational effects, on the surface of the Schwarzschild black hole tending to infinity with a finite volume. The Barrow entropy is parameterized by the index $0\leq\Delta\leq 1$, which depicts the corresponding quantum gravity effects. After Barrow’s discovery, a broad range of papers have been dedicated to study the gravitational and cosmological situations from the Barrow entropy perspective such as Barrow holographic dark energy (BHDE) models Saridakis:2020zol ; Dabrowski:2020atl ; Huang:2021zgj ; Nojiri:2021jxf ; Luciano:2022viz ; Chanda:2022tpk , inflation theory Maity:2022gdy ; Luciano:2023roh ; Saha:2024dhr and black holes physics Abreu:2020dyu ; Abreu:2020wbz ; Abreu:2021kwu ; Abreu:2022pil ; Jawad:2022lww ; Wang:2022hun ; Abreu:2022pil .

### v2

@@ I Introduction
Black holes provide a framework to pursue an essential link between gravitational physics and thermodynamics Bardeen:1973gs . Moreover, further studies by Bekenstein and Hawking confirmed that quantum physics plays a noteworthy role in analyzing the thermodynamic quantities of black holes Bekenstein:1973ur ; Hawking:1975vcx . In 1973, Bekenstein claimed that the black hole entropy, as a thermodynamic quantity, directly connects to its horizon area as a geometrical object Bekenstein:1972tm ; Bekenstein:1973ur . Later, Hawking approved Bekenstein’s conjecture by introducing the Bekenstein–Hawking (BH) formula $S=A/4$. From a purely classical thermodynamics perspective, black holes act like a perfect absorber, without any thermal emission, with an absolute zero temperature. However, Hawking discovered black holes emit thermal radiation in the presence of quantum effects Hawking:1974rv ; Hawking:1975vcx . Hence, black holes can be understood in terms of the generalized laws of thermodynamics with quantum effects Bekenstein:1972tm ; Bekenstein:1973ur ; Hawking:1974rv ; Hod:2015iig . In principle, the mentioned quantum effects allow us to present an interpretation of the laws of black hole mechanics as physically corresponding to the ordinary laws of thermodynamics Wald:1999vt ; Page:2004xp ; Carlip:2014pma ; Almheiri:2020cfm . This point can be seen when we deal with the universal bounds on the entropy of black hole systems in the context of the generalized second law of thermodynamics, telling us the entropy of a closed physical system will grow until it reaches its maximum value. The first attempt to limit the process of entropy growth of a physical system was made by Bekenstein using heuristic arguments of black holes in order to find an upper bound for the entropy of macroscopic bodies in terms of the total energy and radius of the smallest sphere that encloses the object Bekenstein:1980jp . Although the Bekenstein bound conjecture was obtained by studying the physics of black holes, it does not contain Newton’s constant $G$. Therefore, the Bekenstein bound is not restricted only to gravitational scenarios. Despite the quantum proofs of Bekenstein’s discovery Bekenstein:1981zz ; Bekenstein:1982zv ; Bekenstein:2000ai ; Schiffer:1989et ; Bekenstein:1990du ; Casini:2008cr , several counterexamples point out its possible shortcomings Unruh:1982ic ; Unwin:1982pg ; Page:1982fj ; Bekenstein:1983iq ; Unruh:1983ir ; Pelath:1999xt . Further studies by ’t Hooft and Susskind showed that the Bekenstein bound can not be applied to thermodynamic systems involving strong gravity tHooft:1993dmi ; Susskind:1994vu . They claimed that the entropy of our gravitational system should be bounded by area in the context of the holographic principle. Subsequently, Bousso argued that such a spacelike entropy bound is violated more broadly in many dynamical and cosmological situations. He suggested the covariant entropy bound or Bousso entropy bound by introducing a "lightsheet", defined as a region traced out by non-expanding light-rays emitted orthogonally from an arbitrary surface Bousso:1999xy . A derivation of the Bousso bound was given by Flanagan, Marolf, and Wald by considering a condition on a null hypersurface Flanagan:1999jp .
@@
In addition to the BH statistics, we might work with non-Gaussian statistics by taking into account some corrections to the BH entropy coming from non-extensivity of the horizon degrees of freedom (DOF) or quantum gravitational effects on the black hole surface area. These statistics provide a natural framework to study semiclassical effects in the context of the generalized uncertainty principle (GUP). For instance, Tsallis proposed a class of non-Gaussian entropies by considering some non-extensive effects to the Boltzmann-Gibbs statistics Tsallis:1987eu , which is parameterized by index $q$. The non-extensivity parameter $q$ refers to super-extensivity and sub-extensivity when $q<1$ and $q>1$, respectively Tsallis:2002tp . In fact, rare events occur for $10\leq q<$ 1 and frequent events occur for $q>1$ Tsallis:1998as ; NIVEN2004444 , pointing to the stretching or compressing of the entropy curve to lower or higher maximum entropy positions. The consequences of these non-extensivity effects on the thermodynamic parameters have been discussed in Abe:2001ss . In the context of the "gravity-thermodynamics" conjecture, a broad range of gravitational and cosmological scenarios have been studied in the presence of non-extensivity. Such as dark energy (DE) Barboza:2014yfe ; Nunes:2015xsa ; Zadeh:2018poj ; Ghaffari:2018wks ; Tavayef:2018xwx ; SayahianJahromi:2018irq ; Saridakis:2018unr ; Sheykhi:2018dpn ; Lymperis:2018iuz ; Sheykhi:2019bsh ; Huang:2019hex ; Aditya:2019bbk ; DAgostino:2019wko ; Mamon:2020wnh ; Mohammadi:2021wde ; Dheepika:2021fqv ; Nojiri:2022dkr ; Nojiri:2022aof ; Nojiri:2022dkr , cosmic inflation Keskin:2023ngx ; Odintsov:2023vpj ; Teimoori:2023hpv , black holes deOliveira:2005jy ; Tsallis:2012js ; Komatsu:2013qia ; Mejrhit:2020dpo ; Abreu:2020wbz ; Nojiri:2021czz ; Nojiri:2023ikl , gravitational waves Jizba:2024klq , the cosmic microwave background (CMB) Bernui:2005hq ; Bernui:2007wj and modified Newtonian dynamics (MOND) Abreu:2014dna ; Abreu:2018pua . A relativistic generalization of the Boltzmann-Gibbs statistics was introduced by Kaniadakis Kana , which naturally emerges from the special relativity perspective Kaniadakis:2002zz . In $\kappa$-statistics, we deal with a non-exponential distribution function parameterized by the index $-1<\kappa<1$. See its applications in holographic dark energy (HDE) Drepanou:2021jiv ; Hernandez-Almada:2021aiw ; P:2022amn , cosmic inflation Nojiri:2022dkr ; Odintsov:2023vpj ; Lambiase:2023ryq , black holes Abreu:2021avp ; Abreu:2021kwu ; Cimidiker:2023kle and other cosmological situations Luciano:2022eio . Another class of non-Gaussian statistics deals with logarithmic and power-law correction terms due to quantum gravitational effects in the physics of black holes. Considering self-gravitation and backreaction effects to the metric of the black hole navigates us to the BH logarithmic-corrected statistics described by two dimensionless and model-dependent constants $\alpha$ and $\beta$. For gravitational and cosmological applications of the logarithmic-corrected version of the BH entropy, see Refs.Jamil:2010xq ; Cai:2010zw ; Abreu:2020wbz ; Abreu:2020dyu . Recently, Barrow Barrow:2020tzx has proposed a new type of statistics by considering a fractal structure, coming from quantum gravitational effects, on the surface of the Schwarzschild black hole tending to infinity with a finite volume. The Barrow entropy is parameterized by the index $0\leq\Delta\leq 1$, which depicts the corresponding quantum gravity effects. Note that the quantum gravity effects considered in the Barrow entropy are different than the assumed quantum corrections in the context of the logarithmic-corrected entropy Kaul:2000kf , however, it resembles Tsallis entropy Tsallis:1987eu , nevertheless the involved foundations and physical principles are completely different. Also, the discussed fractal behavior does not arise from specific quantum gravity calculations, but from general simple physical principles, which adds to its plausibility and hence it is valid as a first approach to the subject Barrow:2020tzx . Besides phenomenological supports, the Barrow entropy is secured by some important theoretical proofs. In Saridakis:2020zol , the authors used the standard holographic principle at a cosmological structure and Barrow entropy. In Saridakis:2020cqq , the authors calculated the entropy time-variation in a universe filled with matter and dark energy fluids, as well as the corresponding quantity for the apparent horizon. They showed that the generalized second law of thermodynamics is valid in the context of the Barrow statistics due to a small deviation from the Bekenstein-Hawking entropy ($\Delta=0$). Also, in Saridakis:2020lrg , the authors discussed modified cosmological scenarios using Barrow entropy. Moreover, Barrow’s entropy has been tested, theoretically and observationally, by studying the gravitational and cosmological situations such as Barrow holographic dark energy (BHDE) models Dabrowski:2020atl ; Huang:2021zgj ; Nojiri:2021jxf ; Luciano:2022viz ; Chanda:2022tpk , inflation theory Maity:2022gdy ; Luciano:2023roh ; Saha:2024dhr and black holes physics Abreu:2020dyu ; Abreu:2020wbz ; Abreu:2021kwu ; Abreu:2022pil ; Jawad:2022lww ; Wang:2022hun ; Abreu:2022pil ; DiGennaro:2022grw ; Nojiri:2022sfd .
@@ III Conclusions
In summary, the standard Bekenstein bound is partially satisfied in non-Gaussian statistics since its upper limit (i.e. $R_{\text{B}}=1$ for Barrow, $R_{\text{T}}=1$ for Tsallis and $R_{\text{K}}=1$ for Kaniadakis) is not achievable, while the GUP version of the Bekenstein bound provides a framework for a fully satisfied black hole bound in non-Gaussian statistics due to the increase in entropy, which usually happens for $\beta<0$.